Italian Queue 2014 Prev Next

Jun 17-18

Observer: E. Sani, A. Marchetti, M. Fumana
Telescope Operator: D. Gonzales-Huerta
Support Astronomer: D. Thompson on call
Instrument: MODS/LUCI

Summary

At the beginning of the night we are closed due to heavy clouds with chance of rain and strong wind gusts.
When we open the dome the sky is only partly cloudy, the seeing is either bad and gets better only at the end of the night when the humidity increases.
We had condensation on LUCI during all the night.
Guide prove error.
Flexure compensation error
MOS Acquisition Gui got stuck.

Data Summary

Project_RUN Objects Filter ExpTime Notes
Beltran_01 A4 H2 32m out of specs, humidity
Magrini_24 Mask1 210_zJHK, H 50min out of specs, no guiding

Details

* We run dark calibrations for the Magrini_24 program luci.201406180001-0010

05.20 the TO opens the Dome
  • still some wind gusts but the sky is now clear

05.54 We slew to Beltran_01 A4
  • executeLUCIscript.sh A4_H2_DJT.img luci.20140618.0011-0022
  • the average seeing on the GCS is 1.3 at the beginning of the integration, with clear sky.
  • We got an error from the flexure compensation. We switched it off and than on and the problem was fixed, but image luci.20140618.0011 was taken with flexure compensation off.

06.28 DOME CLOSED due to incoming clouds

07:37 TO opens the Dome

07:38 We slew to Beltran_01 A4
  • executeLUCIscript.sh A4_H2_DJT.img luci.20140618.0023-0025
  • at the beginning of the integration the seeing is 1.5 on the GCS

07.49 DOME closed again due to unexpected strong wind, also the temperature is decreasing and is reaching the dewpoint

08:46 TO opens the Dome

08:55 We slew to Cresci campoB_Ks.img
  • target is too faint due to a clouds, we have to change target

08:57 We slew to Magrini

09:07 executeLUCIscript.sh 7C1756Mask1_MOS_210_H_tell.spec for Magrini telluric
  • before acquisition: two files wrong because we did not acquire the target

09:09 executeLUCIscript.sh 7C1756Mask1_MOS_210_HeIJ_tell.acq luci.20140618.00??-0034
  • executeLUCIscript.sh 7C1756Mask1_MOS_210_H_tell.spec luci.20140618.0035-0040

09:17 executeLUCIscript.sh 7C1756Mask1_MOS_210_H.acq luci.20140618.0041-0044
  • The MOS Acquisition gui got stuck. We need to call the SA that suggest to exit the NX session. It worked
  • Given the restart of NX system the last acquisition image (though slit) is actually named luci.20140618..fits instead of 0044. We thus insert by hand the 0045 counting in the "New readout setup".

09.52 executeLUCIscript.sh luci.20140618.0045-0054
  • Error from the Probe because we stayed too long in the same position. The SA suggest to give offsets by hand (yesterday happened something similar, but with the acquisition crash). The recovery takes a while because we need to be sure to go back in the initial position. It took us more 3 iterations.
  • taken during twilight, mostly to be repeated.

11:23 We slew to a blank field to fermorm sky flat for the Beltran_01 program in the H2 filter.
  • We set manually new values on the readout control, but got Error 6. We have to call the SA that suggests to kill the ReadoutManagergGUI and then type
  • observer% cd ./bin
  • observer% ./start_RMGUI.sh from the nx terminal. Did not work.
  • Try again bui exiting the GEIRS terminal and killing the ReadoutManagerGUI, then type on two different nx terminals and from the ./bin directory:
  • observer% ./clean_geirs.sh
  • observer% ./start_geirs.sh and from the over observer% ./start_RMGUI.sh
  • the sky is too bright to continue scientific operations.

11:40 The TO closes the dome
  • We run modsSleeep for MODS and exit the nx client for LUCI.
Topic revision: r8 - 23 Jun 2014, EleonoraSani
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback